# Equi-RO: A 4D mmWave Radar Odometry via Equivariant Networks

### v1

@@ IV-D Ablation Study
We also visualize the trajectories generated by each ablated variant across different splits in Fig. 9. These trajectories demonstrate that the full Equi-RO model consistently achieves the best performance. All three ablated variants show notable performance degradation, particularly in scenarios involving large rotations. The NEN variant, which discards equivariant network design, exhibits the worst results across most splits, underscoring the critical importance of equivariant architecture in our framework.
@@ IV-E Limitations
Although Equi-RO demonstrates superior accuracy and robustness, its performance is affected by the relatively low resolution of radars, as shown in the self-collected dataset. Additionally, while our approach meets general real-time requirements with a processing speed of 76.75 ms per frame on a single NVIDIA A100 GPU, further algorithmic optimization is necessary for deployment on on-board platforms with stricter latency budgets.
@@ V Conclusion
In this paper, we propose a novel 4D mmWave radar odometry algorithm based on equivariant networks. Our method effectively handles noisy, sparse radar point clouds while preserving geometric consistency under large rotations. Extensive experiments on NTU4DRadLM and the self-collected dataset validate its effectiveness and generalizability. This work highlights the potential of equivariant network design for radar-based perception and provides a new perspective for point cloud processing in robotics and autonomous systems. Future work will extend the framework to multi-modal fusion with IMUs and cameras and explore its application to other 3D point clouds with physical attributes.

### v2

@@ IV-D Ablation Study
We also visualize the trajectories generated by each ablated variant across different splits in Fig. 9. These trajectories demonstrate that the full Equi-RO model consistently achieves the best performance. All five ablated variants show notable performance degradation. The variant II, which discards equivariant network design, exhibits the worst results across most splits, underscoring the critical importance of equivariant architecture in our framework.
As shown in Fig. 10, we compare the per-frame average translation and rotation errors of the Equi-RO method with the no invariant velocity compensation variant I across different dynamic-point ratios. As the ratio of dynamic points increases, the errors of the variant I consistently grow, while the full method remains relatively stable. This demonstrates that compensated invariant Doppler features effectively improve robustness in dynamic scenes.
@@ IV-E Limitations
Although Equi-RO demonstrates superior accuracy and robustness, its performance is still affected by the relatively low resolution of radars, as shown in the self-collected dataset. Moreover, while our method achieves real-time performance at 58.15 ms per frame on a single NVIDIA A100 GPU, further optimization is necessary for deployment on on-board platforms. Efficiency can be improved via C++ migration and TensorRT acceleration, bridging the gap between the research prototype and mobile platform constraints.
@@ V Conclusion
In this paper, we propose a novel 4D mmWave radar odometry algorithm based on equivariant networks. Our method effectively handles noisy, sparse radar point clouds while preserving geometric consistency under large rotations. Extensive experiments on NTU4DRadLM and the self-collected dataset validate its effectiveness and generalizability. This work highlights the potential of equivariant network design for radar-based perception and provides a new perspective for point cloud processing in robotics and autonomous systems. Future work will extend the framework to multi-modal fusion with IMUs and cameras, and explore its application to other 3D point clouds with physical attributes.